5kg : 7.5kg = Rs 7.50 : Rs 5
 5kg : 7.5kg = Rs 7.50 : Rs 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
The given statement is False.
Consider the two ratios 5kg: 7.5kg = Rs 7.50: Rs 5 5/7.5 = 7.50/5 50/75
= 75/50

50/75 is further simplified by dividing both numerator and denominator by 25
we get, = 2/3
75/50 is further simplified by dividing both numerator and denominator by 25
we get, = 3/2

Hence, 2/3≠3/2
Therefore, 5kg: 7.5kg≠Rs 7.50: Rs 5


Thanks